the mass m of oxygen in a steel glass cylinder varies directly with the pressure P and inversely with Kelvin temperature T. If the cylinder holds 28 grams of oxygen when the pressure is 4.5 atmospheres and the temperature is 298 K, how much oxygen is in the cylinder if the pressure is 2.8 atmospheres and the temperature is 236 K? round to the nearest tenth of a gram.
A. 28 g
B. 22. g
C. 23.1 g
D. 22.6 g

Answer by mananth(16949)

m= pk/t
28=4.5/298 *k
k=28*298/4.5
k=1854.22
m= 1854.22*2.8/236
m= 21.999
=22g